OpenStudy (anonymous):

Determine whether the relationship represents a positive correlation, negative correlation or no correlation. The amount of interest earned on an investment as time increases. Answer positive negative no relation

OpenStudy (anonymous):

@primeralph

OpenStudy (primeralph):

Positive.

OpenStudy (primeralph):

Got it?

OpenStudy (anonymous):

yes thank you <3

OpenStudy (anonymous):

Write an equation to define the following scenario. You went to an amusement park, paying $12.75 to get in. After that you paid $3 per ride. Answer y = 3x + 12.75 y = 12.75x + 3 y = 3x – 12.75 y = 12.75x – 3

OpenStudy (anonymous):

@primeralph

OpenStudy (primeralph):

A.

OpenStudy (anonymous):

State whether the sequence is arithmetic. If so, state the common difference. 1/4, 1/2, 1, ... Answer 1/2 -1/2 not arithmetic 1 1/2

OpenStudy (primeralph):

Not.

OpenStudy (anonymous):

Name the next two numbers in the pattern 100, 50, 25, 12.5, ... Answer 0, -12.5 6.25, 3.125 62.5, 31.25 0, 12.5

OpenStudy (anonymous):

@primeralph